OpenStudy (anonymous):

multiply x^1/2 * x^3/7

OpenStudy (tyteen4a03):

According to the laws of indices, a^m * a^n = a*(m+n). You're looking at x^(1/2+3/7).

OpenStudy (anonymous):

so would it be 6/14+7/14= 13/14? @tyteen4a03

OpenStudy (nubeer):

yes.. it would be in power of x

OpenStudy (anonymous):

Thank you! @nubeer and @tyteen4a03

OpenStudy (nubeer):

you're welcome ..